# Towards Goal-oriented Prompt Engineering for Large Language Models: A Survey

### v2

@@ Appendix A Peroformance Comparison
Reasoning Arithmetic Reasoning ①②③④⑤ CoT Wei et al. (2022), Self-consistency Wang et al. (2022), Least-to-most Zhou et al. (2022), PoT Chen et al. (2022), Self-refine Madaan et al. (2023), Plan-and-solve Wang et al. (2023b), MWP Zhang et al. (2023), RAP Hao et al. (2023), SelfCheck Miao et al. (2023), Toolformer Schick et al. (2023), Successive Prompt Dua et al. (2022)
Commonsense Reasoning CoT Wei et al. (2022), Self-consistency Wang et al. (2022), MCR Yoran et al. (2023), Plan-and-solve Wang et al. (2023b), Toolformer Schick et al. (2023)
Symbolic Reasoning Self-consistency Wang et al. (2022), Plan-and-solve Wang et al. (2023b)
Logical Reasoning Selection-Inference Creswell et al. (2022), DecomP Khot et al. (2022), RAP Hao et al. (2023)
Planning Virtual Environment ①②④ Zero-shot Planner Huang et al. (2022a), Re-prompting Raman et al. (2022), SALP Gramopadhye and Szafir (2023), RAP Hao et al. (2023), Reflexion Shinn et al. (2023), LLM+P Liu et al. (2023a), SayPlan Rana et al. (2023), DEPS Wang et al. (2023d), GITM Zhu et al. (2023), Inner Monologue Huang et al. (2022b)
Real Environment Inner Monologue Huang et al. (2022b)
Question Answering Multihop QA ①③④⑤ DecomP Khot et al. (2022), Self-ask Press et al. (2022), Reflexion Shinn et al. (2023), MCR Yoran et al. (2023), PEARL Sun et al. (2023)
@@
Dialogue ①④⑤ ProCoT Deng et al. (2023), GDP-Zero Yu et al. (2023), Self-refine Madaan et al. (2023)
Recommendation ③⑤ Recmind Wang et al. (2023c)
Table 9: Applicable tasks from existing works. The stage column indicates the stages involved in the corresponding methods with regard to the goal-oriented framework in Fig. 1.

### v3

@@ Appendix B Performance Comparison
Reasoning Arithmetic CoT Wei et al. (2022), Least-to-most Zhou et al. (2022), Plan-and-solve Wang et al. (2023d), Successive Prompt Dua et al. (2022), RoT Lee and Kim (2023)
MWP Zhang et al. (2023)
PoT Chen et al. (2022), Toolformer Schick et al. (2023), Faithful CoT Lyu et al. (2023), PAL Gao et al. (2023), MathPrompter Imani et al. (2023)
Self-refine Madaan et al. (2023), RAP Hao et al. (2023), SelfCheck Miao et al. (2023), REFINER Paul et al. (2024), CRITIC Gou et al. (2023), MAF Nathani et al. (2023)
Self-consistency Wang et al. (2022)
Commonsense CoT Wei et al. (2022), Plan-and-solve Wang et al. (2023d)
Toolformer Schick et al. (2023)
Self-consistency Wang et al. (2022), MCR Yoran et al. (2023)
Symbolic Plan-and-solve Wang et al. (2023d)
PAL Gao et al. (2023)
Self-consistency Wang et al. (2022)
Logical DecomP Khot et al. (2022)
LINC Olausson et al. (2023), Logic-LM Pan et al. (2023)
RAP Hao et al. (2023)
Selection-Inference Creswell et al. (2022)
Planning Virtual Env LLM+P Liu et al. (2023a), SayPlan Rana et al. (2023), DEPS Wang et al. (2023f), GITM Zhu et al. (2023)
Zero-shot Planner Huang et al. (2022b), Re-prompting Raman et al. (2022), SALP Gramopadhye and Szafir (2023), SayPlan Rana et al. (2023),
Re-prompting Raman et al. (2022), RAP Hao et al. (2023), Reflexion Shinn et al. (2023), SayPlan Rana et al. (2023), DEPS Wang et al. (2023f), GITM Zhu et al. (2023), Inner Monologue Huang et al. (2022c), LLM-Planner Song et al. (2023)
Real Env Inner Monologue Huang et al. (2022c)
Question Answering Multihop QA DecomP Khot et al. (2022), Self-ask Press et al. (2022a)
PEARL Sun et al. (2023)
Faithful CoT Lyu et al. (2023)
Reflexion Shinn et al. (2023), CRITIC Gou et al. (2023), Verify-and-Edit Zhao et al. (2023)
MCR Yoran et al. (2023)
@@
Dialogue ProCoT Deng et al. (2023), RLP Fischer (2023), Cue-CoT Wang et al. (2023c), SAFARI Wang et al. (2023b)
Self-refine Madaan et al. (2023)
GDP-Zero Yu et al. (2023a),
Recommendation Recmind Wang et al. (2023e), DOKE Yao et al. (2023a), InteRecAgent Huang et al. (2023b)
Recmind Wang et al. (2023e)
Table 10: Applicable tasks from existing works. The stage column indicates the stages involved in the corresponding methods with regard to the goal-oriented framework in Fig. 1. Note that the methods in the table are not exclusively applicable to the tasks listed in the table to which they belong. We list them based on the evaluated tasks in their original papers.